What is escape velocity
1
Expert's answer
2021-01-01T14:05:13-0500

Answer

According to physics theory velocity which is required to escape from earth 'science gravitational field.

Which represented as below

"v_e=\sqrt{\frac{2GM}{R}}=\sqrt{2gR}" =11.2m/s